# Revisiting End-to-End Learning with Slide-level Supervision in Computational Pathology

### v1

@@ Abstract
Pre-trained encoders for offline feature extraction followed by multiple instance learning (MIL) aggregators have become the dominant paradigm in computational pathology (CPath), benefiting cancer diagnosis and prognosis. However, performance limitations arise from the absence of encoder fine-tuning for downstream tasks and disjoint optimization with MIL. While slide-level supervised end-to-end (E2E) learning is an intuitive solution to this issue, it faces challenges such as high computational demands and suboptimal results. These limitations motivate us to revisit E2E learning. We argue that prior work neglects inherent E2E optimization challenges, leading to performance disparities compared to traditional two-stage methods. In this paper, we pioneer the elucidation of optimization challenge caused by sparse-attention MIL and propose a novel MIL called ABMILX. It mitigates this problem through global correlation-based attention refinement and multi-head mechanisms. With the efficient multi-scale random patch sampling strategy, an E2E trained ResNet with ABMILX surpasses SOTA foundation models under the two-stage paradigm across multiple challenging benchmarks, while remaining computationally efficient ($<$ 10 RTX3090 hours). We show the potential of E2E learning in CPath and calls for greater research focus in this area. The code is here.
Figure 1: (a,b) We compare end-to-end trained ResNet with various foundation models using two-stage paradigm in terms of performance, model size, and pretraining data. This shows the performance potential of end-to-end learning for CPath under low computational budget. (c) Compared to sampling strategies, different MILs have a more significant impact and lower cost on E2E learning.
@@ 1 Introduction
To address these challenges, Lu et al. [38] proposed a two-stage paradigm based on multiple instance learning (MIL) [41], allowing efficient WSI analysis without fine-grained annotations. This approach first divides each WSI (a bag) into thousands of image patches (instances). Pretrained encoders extract offline instance features, which are then aggregated into bag features by a sparse-attention MIL model, ultimately leading to slide prediction. By operating in the latent space rather than images, this paradigm enables slide-level supervised model training within reasonable memory constraints. However, its performance heavily depends on the quality of offline features [9]. To improve offline feature quality, a series of pathology foundation models [10, 58, 24, 60] (FMs) like UNI [10] and GigaPath [60] have been developed. As shown in Figure 1, despite scaling data volume to 170K WSIs ($>$200TB) and model size over 1B, these approaches still perform unsatisfactorily on specific tasks. We attribute this to the lack of unified optimization in the two-stage paradigm, resulting in encoders with insufficient adaptation of downstream task and disjoint optimization with MIL models.
@@
Furthermore, we adopt simple but effective multi-scale random patch sampling to incorporate multi-scale information while reducing E2E learning computational costs. Our E2E learning framework achieves significant performance improvements (e.g., $+20\%$ accuracy on PANDA) while maintaining computationally efficient ($<$ 10 RTX3090 GPU hours on TCGA-BRCA). The main contributions can be summarized as follows:
@@ 2 Related Works
Computational Pathology. The advent of WSI in computational pathology (CPath) has revolutionized approaches to cancer diagnosis and prognosis by furnishing a comprehensive, high-resolution view of tissue specimens [14, 51, 13]. Due to processing gigapixel images is computationally intensive, traditional CPath methods have adopted a two-stage paradigm to prioritize efficiency [38]. In the first stage, an offline feature extractor, typically pre-trained in general datasets [38, 21] or pathology datasets [10, 58, 24, 60], is employed to encode tissue patches into features. In the subsequent stage, MIL are applied to aggregate these features for slide-level prediction. Most research [26, 31, 49, 53, 33, 52, 64, 32, 19] has focused on refining MIL stage with advanced aggregation mechanisms, notably the use of sparse attention. MIL model computes attention scores for each patch and aggregates only the most informative ones [25], thereby reducing noise and enhancing slide-level prediction accuracy in WSIs with scattered key histological features. Some studies have sought to better exploit information contained in WSIs by directly extracting supplementary visual cues from the entire slide [50, 18]. Others have refined the extracted features to better match the dataset through either multi-stage feature extractor fine-tuning [30] or online instance feature re-embedding [54] to more precisely tailor the extracted features to the dataset. However, current two-stage approaches rely on pretrained offline feature extractors that are not jointly optimized with the MIL model, potentially resulting in features that inadequately capture the complex nuances of pathology data in WSIs [30]. In this context, E2E approaches have emerged as a promising paradigm.
@@ 3.1 Slide-level Supervised End-to-End Learning
where $\mathcal{V}_{\mathcal{S}}(\cdot)$ denotes the function of vanilla random sampling. It is notable that we set $\sum_{j=1}^{t}(\sigma_{j})=1$ to ensure that $\sum_{j=1}^{t}(\hat{s}_{j})=s$. We resize the sampled instances of different scales and context extents $\{\hat{\bm{L}}_{1},\cdots,\hat{\bm{L}}_{j},\cdots,\hat{\bm{L}}_{t}\}$ to a unified resolution and merge them as the final sampling set $\bm{L}$. On the one hand, multi-scale sampling simulates the multi-scale perspective of pathologists during diagnosis and improves the CPath performance of our method. On the other hand, the unified resolution for different context avoids the additional cost and maintains parallel training, while remaining the different scale perspectives of original instance. Appendix give more details about sampling.
@@ 3.2 ABMILX for Effective End-to-End Learning
Multi-head Local Attention. Considering that the false attention from under-converged MIL usually exhibit a random distribution, we propose a multi-head local attention module (MHLA) to directly suppress the excessive focus on redundant instances while improving the attention for the discriminative ones. Specifically, we divide the features of all sampled instances $\bm{E}$ into $m$ head features $\{\bm{H}^{1},\cdots.\bm{H}^{j}.\cdots.\bm{H}^{m}\}$, where $\bm{H}^{j}\in\mathbb{R}^{s\times\lceil D/m\rceil}$. Within each head, the head features are input into a shared MLP to compute the corresponding attention, $\bm{A}^{j}=\text{MLP}(\bm{H}^{j})$. The $\bm{A}^{j}\in\mathbb{R}^{s\times 1}$ denotes the local attention vector of the $j$-th head, which possesses sparse characteristic important in CPath tasks , more details seen in Appendix. In the E2E learning, the separate voting from multiple heads allows to reduce the excessive focus on redundant instances, while the attention from different feature subspaces helps to provide a more comprehensive view on discriminative instances. Finally, we aggregate the features within each head through $\bm{A}^{j}$ to obtain the head-level slide features $(\bm{Z}^{1},\cdots,\bm{Z}^{j},\cdots,\bm{Z}^{m})$, which are then concatenated as the final slide feature :
@@
Therefore, ABMILX utilizes the $\bm{A}^{j}$ as prior distribution to grant sparse discriminative instances with higher propagation weights to find more potential instances while suppressing the normal ones. More theoretical analysis about ABMILX is available in Appendix.
@@ 3.3 Sparse Attention Analysis in E2E Learning
Different MILs in E2E Sparsity Sub.$\uparrow$ Surv.$\uparrow$
@@
To intuitively analyze the effect of sparse attention on E2E training, we quantitative the sparsity of different MILs by the proportion of activated patches. Sparsity is statistically derived from the CAMELYON dataset [3]. Moreover, the right figure visualizes attention scores (bottom) and corresponding distribution (middle) of MILs during training. We demonstrate the following: (1) In E2E optimization, extreme sparsity causes ABMIL to overlook discriminative regions while over-focusing on redundant ones, leading to worst performance. (2) Although the global attention of TransMIL eliminates this extreme sparsity and covers some discriminative regions, it is also largely distracted by the redundant ones, which also brings limited accuracy gains. (3) In contrast, both MHLA and A+ make ABMILX maintains reasonably sparse attention, which considers most of the discriminative regions while maintaining low attention to normal patches. Benefited from them, ABMILX achieves the best performance in different C-Path tasks. Besides, learnable $\alpha$ also helps adaptively adjusting the sparsity and brings more accuracy gains. More experiments and analysis about the affect of different MILs in E2E learning are available in Sec. 4.3 and Appendix.
@@ 4.1 Datasets and Evaluation Metrics
We use PANDA [5], TCGA-BRCA, and TCGA-NSCLC to evaluate the performance in cancer grading and sub-typing tasks. For cancer prognosis, we use TCGA-LUAD, TCGA-BRCA, TCGA-BLCA to evaluate performance on the survival analysis task. For cancer grading, we evaluate model performance using top-1 accuracy (Acc.). And area under the ROC curve (AUC) is used for sub-typing. For survival analysis, we employ the concordance index (C-index) [20]. To ensure robust statistical evaluation, we conducted a 1000-time bootstrapping evaluation and report the mean and 95% confidence interval. Please refer to Appendix for more details.
@@
Table 2: Performance comparison across ISUP grading (PANDA) and survival analysis.
Encoder Method E2E PANDA (Acc.$\uparrow$) LUAD BRCA BLCA
@@ 4.2 Main Results
ABMILX Shows E2E Potential. Through E2E learning with ABMILX and downstream data, we achieve FMs-level performance using ResNet models that were pre-trained in ImageNet-1k. It outperforms FMs on multiple challenging datasets (+4% Acc. on PANDA, +0.8% AUC on BRCA). Moreover, the E2E learning cost of ABMILX is substantially lower than the pretraining cost of FMs, approaching the cost of training second-stage aggregators, with more details provided in next section. Additionally, we show that fine-tuning upstream pre-trained aggregators, like CHIEF and GIGAP, did not yield the desired results. This further underscores the necessity of E2E training of encoders and aggregators with slide supervision. In particular, we demonstrate the scalability of the proposed method with respect to the model size. Except for survival analysis influenced by the sampling numbers (Table 2), R50 shows a general improvement compared to R18. In conclusion, empowered by ABMILX, we present the impact and enormous potential of E2E learning in CPath. Appendix presents more discussion about other E2E methods.
@@
4.3 Ablation Study
MIL Matters in End-to-End Trainings. Right Table shows the impact of the sampling and aggregation modules on E2E learning. We observe that different MILs have a significant effect on E2E learning performance. Specifically, ABMIL exhibits unsatisfactory performance in E2E training, except on the PANDA dataset. We attribute this to its excessive sparsity hindering E2E optimization. The PANDA dataset contains fewer patches per slide (500 vs. 10,000 for TCGA-BRCA), enabling MIL to focus on discriminative regions more easily, thus suffering minimal impact on E2E optimization. RRTMIL exacerbates this problem, leading to optimization collapse. This complex MIL method, with a serial feature re-embedding module preceding ABMIL, makes E2E training more fragile. It further impairs the representation of features already affected by sparse attention, accelerating the collapse of the optimization loop. TransMIL and DSMIL, the transformer-based methods, partially mitigate this issue. However, relying solely on global attention struggles to focus on key regions within the numerous redundant patches in training, resulting in a considerable performance gap compared to FMs. ABMILX, while maintaining desirable sparsity, alleviates optimization issues and achieves significant performance improvements. Furthermore, complex sampling strategies, such as attention-based sampling, offer only limited performance gains compared to vanilla random sampling. Such strategies require patch evaluation and incur substantial training time (TTime). Multi-scale random instance sampling (MRIS) shows better performance. Appendix provides further discussion.
@@ 4.3 Ablation Study
CHIEF [58] ABMIL [25] 32GB×$\times$ - h 1+2h 2G 6.2s 65.66 91.09 64.02
UNI [10] TransMIL [46] 80GB×$\times$ - h 1+7h 8G 25s 68.06 93.33 60.45
GIGAP [60] GIGAP [60] 80GB×$\times$3,072h 6+23h 7G 83s 65.86 93.72 62.64
@@
Validity of Our ABMILX. Table 3 (bottom) ablates key components of ABMILX. E2E training with ABMIL performs poorly except for PANDA due to optimization challenges. It performs below SOTA MIL with R50 features and significantly underperforming FMs. After introducing multi-head mechanisms, the extreme focus on redundant instances caused by sparse attention is effectively mitigated, thus achieving consistent improvements. More importantly, by refining attention using global patch correlations in the attention plus module, optimization issues are further alleviated. This improvement helps ABMILX achieve FMs-level performance. Furthermore, the sharp performance degradation when freezing the encoder demonstrates the necessity of E2E learning. We also validate ABMILX under two-stage paradigm in Appendix.
@@ Appendix A Theoretical Analysis of Optimization Risk in End-to-End Training
Definition 1. (Optimization Risk) Let $\mathcal{N}$ denote the set of noisy instances, $\mathcal{D}$ denote discriminative instances, and $O(\hat{a})$ is a measure of the contribution of attention $\hat{a}$ to the final bag feature $Z$. $O(\cdot)$ is monotone increasing. We define the optimization risk $\mathcal{R}$ as the impact of maximum attention value among noisy instances to bag feature:

### v2

@@ Abstract
Pre-trained encoders for offline feature extraction followed by multiple instance learning (MIL) aggregators have become the dominant paradigm in computational pathology (CPath), benefiting cancer diagnosis and prognosis. However, performance limitations arise from the absence of encoder fine-tuning for downstream tasks and disjoint optimization with MIL. While slide-level supervised end-to-end (E2E) learning is an intuitive solution to this issue, it faces challenges such as high computational demands and suboptimal results. These limitations motivate us to revisit E2E learning. We argue that prior work neglects inherent E2E optimization challenges, leading to performance disparities compared to traditional two-stage methods. In this paper, we pioneer the elucidation of optimization challenge caused by sparse-attention MIL and propose a novel MIL called ABMILX. ABMILX mitigates this problem through global correlation-based attention refinement and multi-head mechanisms. With the efficient multi-scale random patch sampling strategy, an E2E trained ResNet with ABMILX surpasses SOTA foundation models under the two-stage paradigm across multiple challenging benchmarks, while remaining computationally efficient (<< 10 RTX3090 GPU hours). We demonstrate the potential of E2E learning in CPath and calls for greater research focus in this area. The code is here.
@@ 1 Introduction
Figure 1: (a,b) We compare E2E trained ResNet with various foundation models using two-stage paradigm in terms of performance, model size, and pretraining data. This demonstrates the performance potential of E2E learning for computational pathology under low computational budget. (c) Compared to sampling strategies, different MILs have a more significant impact and lower cost on E2E learning.
To address these challenges, Campanella et al. [7] proposed a two-stage paradigm based on multiple instance learning (MIL) [41], allowing efficient WSI analysis without fine-grained annotations. This approach first divides each WSI (a bag) into thousands of image patches (instances). Pretrained encoders extract offline instance features, which are then aggregated into bag features by a sparse-attention MIL model, ultimately leading to slide prediction. By operating in the latent space rather than images, this paradigm enables slide-level supervised training within reasonable memory constraints. However, its performance heavily depends on the quality of offline features [10]. To improve offline feature quality, a series of pathology foundation models [11, 62, 24, 64] (FMs) like UNI [11] and GigaPath [64] have been developed. As shown in Figure 1, despite scaling data volume to 170K WSIs (>>200TB) and model size over 1B, these approaches still perform unsatisfactorily on specific tasks. We attribute this to the lack of unified optimization in the two-stage paradigm, resulting in encoders with insufficient adaptation of downstream task and disjoint optimization with MIL models.
@@
Furthermore, we adopt simple but effective multi-scale random patch sampling to incorporate multi-scale information while reducing E2E learning computational costs. Our E2E learning framework achieves significant performance improvements (e.g., $20\%$ accuracy on PANDA) while maintaining computationally efficient (<< 10 RTX3090 GPU hours on TCGA-BRCA). The main contributions can be summarized as follows:
@@ 2 Related Works
Computational Pathology. The advent of WSI in computational pathology (CPath) has revolutionized approaches to cancer diagnosis and prognosis by furnishing a comprehensive, high-resolution view of tissue specimens [15, 53, 14]. Due to processing gigapixel images is computationally intensive, traditional CPath methods have adopted a two-stage paradigm to prioritize efficiency [38]. In the first stage, an offline feature extractor, typically pre-trained in general datasets [38, 21] or pathology datasets [11, 62, 24, 64], is employed to encode tissue patches into features. In the subsequent stage, MIL are applied to aggregate these features for slide-level prediction. Most research [26, 31, 51, 55, 58, 33, 54, 68, 32, 57] has focused on refining MIL stage with advanced aggregation mechanisms, notably the use of sparse attention. MIL model computes attention scores for each patch and aggregates only the most informative ones [25], thereby reducing noise and enhancing slide-level prediction accuracy in WSIs with scattered key histological features. Some studies have sought to better exploit information contained in WSIs by directly extracting supplementary visual cues from the entire slide [52, 19]. Others have refined the extracted features to better match the dataset through either multi-stage feature extractor fine-tuning [30] or online instance feature re-embedding [56] to more precisely tailor the extracted features to the dataset. However, current two-stage approaches rely on pretrained offline feature extractors that are not jointly optimized with the MIL model, potentially resulting in features that inadequately capture the complex nuances of pathology data in WSIs [30]. In this context, E2E approaches have emerged as a promising paradigm.
@@ 3.1 Slide-level Supervised End-to-End Learning
where $\mathcal{V}_{\mathcal{S}}(\cdot)$ denotes the function of vanilla random sampling. It is notable that we set $\sum_{j=1}^{t}(\sigma_{j})=1$ to ensure that $\sum_{j=1}^{t}(\hat{s}_{j})=s$. We resize the sampled instances of different scales and context extents $\{\hat{\bm{L}}_{1},\cdots,\hat{\bm{L}}_{j},\cdots,\hat{\bm{L}}_{t}\}$ to a unified resolution and merge them as the final sampling set $\bm{L}$. On the one hand, multi-scale sampling simulates the multi-scale perspective of pathologists during diagnosis and improves the CPath performance of our method. On the other hand, the unified resolution for different context avoids the additional cost and maintains parallel training, while remaining the different scale perspectives of original instance. Appendix C.4 give more details about sampling.
@@ 3.2 ABMILX for Effective End-to-End Learning
Multi-head Local Attention. Considering that the false attention from under-converged MIL usually exhibit a random distribution, we propose a multi-head local attention module (MHLA) to directly suppress the excessive focus on redundant instances while improving the attention for the discriminative ones. Specifically, we divide the features of all sampled instances $\bm{E}$ into $m$ head features $\{\bm{H}^{1},\cdots.\bm{H}^{j}.\cdots.\bm{H}^{m}\}$, where $\bm{H}^{j}\in\mathbb{R}^{s\times\lceil D/m\rceil}$. Within each head, the head features are input into a shared MLP to compute the corresponding attention, $\bm{A}^{j}=\text{MLP}(\bm{H}^{j})$. The $\bm{A}^{j}\in\mathbb{R}^{s\times 1}$ denotes the local attention vector of the $j$-th head, which possesses sparse characteristic important in CPath tasks. In the E2E learning, the separate voting from multiple heads allows to reduce the excessive focus on redundant instances, while the attention from different feature subspaces helps to provide a more comprehensive view on discriminative instances. Finally, we aggregate the features within each head through $\bm{A}^{j}$ to obtain the head-level slide features $(\bm{Z}^{1},\cdots,\bm{Z}^{j},\cdots,\bm{Z}^{m})$, which are then concatenated as the final slide feature :
@@
Therefore, ABMILX utilizes the $\bm{A}^{j}$ as prior distribution to grant sparse discriminative instances with higher propagation weights to find more potential instances while suppressing the normal ones. More theoretical analysis about ABMILX is available in Appendix A.
@@ 3.3 Sparse Attention Analysis in E2E Learning
[…]
@@
To intuitively analyze the effect of sparse attention on E2E training, we quantitative the sparsity of different MILs by the proportion of activated patches. Sparsity is statistically derived from the CAMELYON dataset [3]. Moreover, the right figure visualizes attention scores (bottom) and corresponding distribution (middle) of MILs during training. We demonstrate the following: (1) In E2E optimization, extreme sparsity causes ABMIL to overlook discriminative regions while over-focusing on redundant ones, leading to worst performance. (2) Although the global attention of TransMIL eliminates this extreme sparsity and covers some discriminative regions, it is also largely distracted by the redundant ones, which also brings limited accuracy gains. (3) In contrast, both MHLA and A+ make ABMILX maintains reasonably sparse attention, which considers most of the discriminative regions while maintaining low attention to normal patches. Benefited from them, ABMILX achieves the best performance in different C-Path tasks. Besides, learnable $\alpha$ also helps adaptively adjusting the sparsity and brings more accuracy gains. More experiments and analysis about the affect of different MILs in E2E learning are available in Sec. 4.3 and Appendix C.2.
@@ 4.1 Datasets and Evaluation Metrics
We use PANDA [6], TCGA-BRCA, and TCGA-NSCLC to evaluate the performance in cancer grading and sub-typing tasks. For cancer prognosis, we use TCGA-LUAD, TCGA-BRCA, TCGA-BLCA to evaluate performance on the survival analysis task. For external validation, we use CPTAC-LUAD, CPTAC-LUSC to evaluate the generalization ability. For cancer grading, we evaluate model performance using top-1 accuracy (Acc.). And area under the ROC curve (AUC) is used for sub-typing. For survival analysis, we employ the concordance index (C-index) [20]. To ensure robust statistical evaluation, we conducted a 1000-time bootstrapping evaluation and report the mean and 95% confidence interval. Please refer to Appendix B for more details.
@@
Table 2: Performance comparison across ISUP grading (PANDA) and survival analysis. OOM denotes Out-of-Memory in 24GB-3090.
[…]
@@ 4.2 Main Results
ABMILX Shows E2E Potential. Through E2E learning with ABMILX and downstream data, we achieve FMs-level performance using ResNet models that were pre-trained in ImageNet-1k. It outperforms FMs on multiple challenging datasets (+4% Acc. on PANDA, +0.8% AUC on BRCA). Moreover, the E2E learning cost of ABMILX is substantially lower than the pretraining cost of FMs, approaching the cost of training second-stage aggregators, with more details provided in next section. Additionally, we show that fine-tuning upstream pre-trained aggregators, like CHIEF and GIGAP, did not yield the desired results. This further underscores the necessity of E2E training of encoders and aggregators with slide supervision. In particular, we demonstrate the scalability of the proposed method with respect to the model size. Except for survival analysis influenced by the sampling numbers (Table 2), R50 shows a general improvement compared to R18. Most critically, we validated the generalization ability through external validation on the CPTAC dataset (Table 3). A ResNet-50 encoder trained on TCGA using our E2E framework not only shows superior generalization but also outperforms UNI, a ViT-L pre-trained on over one billion pathology patches. This result validates that our E2E learning approach fosters robust transferability that can overcome cross-dataset domain shifts, rivaling the benefits of massive-scale pre-training. In conclusion, empowered by ABMILX, we present the impact and enormous potential of E2E learning in CPath. We also present more discussion in Appendix C.1.
Table 3: Performance of different methods on external validation from TCGA to CPTAC datasets.
[…]
ResNet-50 ABMIL [25] 66.42 46.34
TransMIL [48] 74.59 48.24
WIKG [33] 64.04 OOM
UNI [11] ABMIL [25] 83.73 53.59
TransMIL [48] 85.24 51.36
WIKG [33] 83.56 OOM
ResNet-50 ABMILX (ours) 85.19 54.00
4.3 Ablation Study
[…]
@@ 4.3 Ablation Study
MIL Matters in End-to-End Trainings. Right Table shows the impact of the sampling and aggregation modules on E2E learning. We observe that different MILs have a significant effect on E2E learning performance. Specifically, ABMIL exhibits unsatisfactory performance in E2E training, except on the PANDA dataset. We attribute this to its excessive sparsity hindering E2E optimization. The PANDA dataset contains fewer patches per slide (500 vs. 10,000 for TCGA-BRCA), enabling MIL to focus on discriminative regions more easily, thus suffering minimal impact on E2E optimization. RRTMIL exacerbates this problem, leading to optimization collapse. This complex MIL method, with a serial feature re-embedding module preceding ABMIL, makes E2E training more fragile. It further impairs the representation of features already affected by sparse attention, accelerating the collapse of the optimization loop. TransMIL and DSMIL, the transformer-based methods, partially mitigate this issue. However, relying solely on global attention struggles to focus on key regions within the numerous redundant patches in training, resulting in a considerable performance gap compared to FMs. ABMILX, while maintaining desirable sparsity, alleviates optimization issues and achieves significant performance improvements. Furthermore, complex sampling strategies, such as attention-based sampling, offer only limited performance gains compared to vanilla random sampling. Such strategies require patch evaluation and incur substantial training time (TTime). Multi-scale random instance sampling (MRIS) shows better performance. Appendix C.2 provide further discussion.
@@
[…]
@@
Validity of Our ABMILX. Table 4 (bottom) ablates key components of ABMILX. E2E training with ABMIL performs poorly except for PANDA due to optimization challenges. It performs below SOTA MIL with R50 features and significantly underperforming FMs. After introducing multi-head mechanisms, the extreme focus on redundant instances caused by sparse attention is effectively mitigated, thus achieving consistent improvements. More importantly, by refining attention using global patch correlations in the attention plus module, optimization issues are further alleviated. This improvement helps ABMILX achieve FMs-level performance. Furthermore, the sharp performance degradation when freezing the encoder demonstrates the necessity of E2E learning. We also validate ABMILX under two-stage paradigm in Appendix C.3.
@@ 5 Conclusion
Acknowledgements
This work was supported by Shenzhen Science and Technology Program (JCYJ20240813114237048), “Science and Technology Yongjiang 2035“ key technology breakthrough plan project (2024Z120), Chinese government-guided local science and technology development fund projects (scientific and technological achievement transfer and transformation projects) (254Z0102G).
@@ Appendix A Theoretical Analysis of Optimization Risk in End-to-End Training
A.1 Definition of Optimization Risk
Definition 1. (Optimization Risk) Let $\mathcal{N}$ denote the set of noisy instances, $\mathcal{D}$ denote discriminative instances, and O(a^)O($\hat${a}) is a measure of the contribution of attention $\hat{a}$ to the final bag feature $Z$. $O(\cdot)$ is monotone increasing. We define the optimization risk $\mathcal{R}$ as the impact of maximum attention value among noisy instances to bag feature: